OpenStudy (anonymous):

Simplify this :

OpenStudy (anonymous):

\[\frac{ [1-(1+3h+3h^2+3h^3)]-(2) }{ h }\]

OpenStudy (lgbasallote):

precalculus? or no?

OpenStudy (anonymous):

yes it is

OpenStudy (lgbasallote):

you should do 1 - (1 + 3h + 3h^2 + 3h^3) first

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

do i just get rid of the first round bracket?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (cruffo):

correct. Distribute the - through [1 -1*( everything inside parenthesis) ]

OpenStudy (anonymous):

oh so you mutiply everything by -1 ?:S

OpenStudy (anonymous):

@cruffo

OpenStudy (cruffo):

sorry had to step away..

OpenStudy (anonymous):

it's alright :P

OpenStudy (cruffo):

correct: \[\frac{ [1-(1+3h+3h^2+3h^3)]-(2) }{ h } = \frac{ [1-1-3h-3h^2-3h^3]-2 }{ h }\]

OpenStudy (cruffo):

but I'm curious: did you originally have to do 1-(1+h)^3 in the top?

OpenStudy (anonymous):

well i have this equation: \[y=1-x^3\] and i need to find the average slope of the line passig through A (-1,2) and Δx=h

OpenStudy (anonymous):

so that was what i came up with

OpenStudy (cruffo):

I think before you go any further, there is something that need to be corrected.

OpenStudy (anonymous):

ohh what ?

OpenStudy (cruffo):

Here x = -1, so you want \[f(-1+h) = 1-(-1+h)^3 = 1-[(-1)^3 + 3(-1)^2h + 3(-1)h^2 + h^3\] \[ = 1-(-1+3h-3h^2+h^3)\] So the top would be \[\frac{ [1-(-1+3h-3h^2+h^3)]-(2) }{ h } = \frac{ [1+1-3h+3h^2-h^3]-2 }{ h }\]

OpenStudy (anonymous):

then the -2 would cancel out with 2

OpenStudy (anonymous):

and we would factor out the h ?

OpenStudy (cruffo):

you got it.

OpenStudy (anonymous):

so m= \[3+3h-h^2\]

OpenStudy (cruffo):

yep.

OpenStudy (anonymous):

oh wrong signs

OpenStudy (cruffo):

hah. right. those pesky negatives...

OpenStudy (anonymous):

but my text book says the final answer is: \[-h2+3h-3\] :S :S

OpenStudy (cruffo):

just rearrange the terms in decending order in h

OpenStudy (anonymous):

ohh it's the same thing .. :$ LOL i thought it was different haha :$

OpenStudy (anonymous):

thank-you so much :D !

OpenStudy (cruffo):

np!

OpenStudy (anonymous):

if part b says to calculate slope for h= 10,2,1,1/2 do i find the limits for his @cruffo ?

OpenStudy (cruffo):

Sorry for the delay.

OpenStudy (cruffo):

Should just be able to plug in the numbers for h in you previous answer, \[\large -h^2 + 3h - 3\]

OpenStudy (anonymous):

thanks :D